OpenStudy (anonymous):

A rectangle has a length of the fifth root of 16 inches and a width of 2 to the 1 over 5 power inches. Find the area of the rectangle. 2 to the 3 over 5 power inches squared 2 to the 4 over 5 power inches squared 2 inches squared 2 to the 2 over 5 power inches squared

OpenStudy (cwrw238):

converting to fractional powers fifth root 0f 16 = 16^(1/5) = (2^4)^(1/5) = 2^(4/5)

OpenStudy (cwrw238):